OpenStudy (my4babiesandme):

state the value of the discriminant and describe the nature of the solutions 14^2+2x+8=0

jimthompson5910 (jim_thompson5910):

D=b^2-4ac D=(2)^2-4(14)(8) D=4-4(14)(8) D=4-448 D=-444 Since the discriminant is less than zero, this means that there are two complex solutions. In other words, there are no real solutions.

OpenStudy (my4babiesandme):

so it would be -444 and the solutions would be imaginary?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (my4babiesandme):

thank you